If possible, completely factor the expression. (if the expression is not factorable, enter NF.)
x^2 - 8x - 9
1
Expert's answer
2013-01-23T09:36:26-0500
x^2 - 8x - 9 = x^2 + x - x - 8x - 9 = x^2+ x -9x - 9 =
=x(x+1) - 9(x+1) = (x-9)*(x+1)
Answer: (x-9)*(x+1)